25. August 2009 — Dienstag

Admin

Backup
  • Manually clean tape drive, EMC 23189374… still broken since 14. April 2008!
  • EMC 31051592:
    • kill off the nsrck that started and hung at midnight.
    • Discover the “undef” volume problem is a side effect of this daily process not working.
    • Continue manual nsrck -L6’s against all clients in /nsr/index
  • EMC 31099632 — open, sev 2 — med104 failing VSS SYSTEM SERVICES:\ backup due to SQL
  • Weekly cloning script finished. So, once the nightly backups finally finish, I’ll be able to recycle networker on the server to see if that clears up the issue with 31051592.

Comments are closed.